OpenStudy (anonymous):

solve the equation; there will be three solutions. 27x^3-64=0

OpenStudy (callisto):

27x^3-64=0 (3x-4)(9x^2 + 12x + 16) =0 (3x-4) =0 or (9x^2 + 12x + 16) =0 Can you solve it from here?

OpenStudy (anonymous):

do you just solve for x now?

OpenStudy (callisto):

Far to my knowledge, you cannot factorise (9x^2 + 12x + 16) For that you need to use quadratics formula

OpenStudy (anonymous):

I concur, I calculated 12x as 24x in my head...

OpenStudy (anonymous):

x^3 = 64/27 -> x= 4/3...

OpenStudy (callisto):

Silly me.... I think too much :S

OpenStudy (anonymous):

Still a problem though....

OpenStudy (callisto):

Best answer is from @estudier :)

OpenStudy (anonymous):

ok i understand where the 4/3 comes from but not the 64/27

OpenStudy (anonymous):

27x^3-64=0

OpenStudy (anonymous):

Ho hum, 3 solutions.....

OpenStudy (callisto):

From (9x^2 + 12x + 16) =0, you can get 2 unreal solutions

OpenStudy (anonymous):

64/27 comes from sending everything to the right : 27x^3-64 = 0 27x^3 = 64 x^3 = 64/27

OpenStudy (anonymous):

OK, so you got 1 real from me and 2 complex from callisto, all done

OpenStudy (callisto):

(9x^2 + 12x + 16) =0 \[x=\frac{-12\pm \sqrt {12^2 -4(9)(16)}}{2(9)} = \frac{-12\pm \sqrt {-432}}{2(9)} =\frac{12(-1\pm \sqrt {-3})}{18}\]\[=\frac{2(-1\pm \sqrt {-3})}{3}\]